# A Szegö limit theorem for a class of Toeplitz operators on the Bergman space of the unit ball with singular symbols Daniel Ivan Ramirez Montaño

### v1

@@ 3 The trace of a multiple composition
For the special case when $A=T_{a\mathrm{d}\sigma}$ is a Toeplitz operator, Fubini’s theorem and the reproducing property of the Bergman kernels yield in the expression
@@
We present a formula for the trace of a composition of $m$ Toeplitz operators associated with the same manifold $\Gamma\subset\mathbb{B}_{n}$.
@@
Let $m\geq 2$ be a natural number, $a_{1},{\scriptstyle\dots},a_{m}\in L^{\infty}(\Gamma)$ with compact support, and ${S:=T_{a_{1}\mathrm{d}\sigma}\circ{\scriptstyle\cdots}\circ T_{a_{m}\mathrm{d}\sigma}}$. Then
@@
Notice that $\frac{1-\langle\xi_{j},\xi_{j+1}\rangle}{1-|\xi_{j}|^{2}}\in(-\infty,0]$ if and only if $\langle\xi_{j},\xi_{j+1}\rangle\in[1,\infty)$. However, this does not occur since $\xi_{j},\xi_{j+1}\in\mathbb{B}_{n}$. It follows from (16) that $\Phi$ is well defined and continuous on $\Gamma^{m}$. Moreover, from (15) we find that all the $\xi_{j}$ involved in the integration are the ones in the support of $a_{j}$. Since each $a_{j}$ has compact support in $\Gamma$, we conclude that there is an open neighborhood of $(-\infty,0]$ such that only values outside this neighborhood are used to evaluate $\operatorname{Log}$ in the expression for $\Phi$ when computing (15).
We describe the behavior of $\mathrm{Im}\big{(}\Phi(\xi_{1},{\scriptstyle\dots},\xi_{m})\big{)}$ with the following result.
@@
4 Asymptotic of the trace of a multiple composition
This section is devoted to finding the asymptotic behavior of the integral expression in (13) when $\alpha$ tends to infinity. Throughout the section, we will keep $m\geq 2$ and $a_{1},{\scriptstyle\dots},a_{m}\in C^{2}_{0}(\Gamma)$ fixed and denote $S:=T_{a_{1}\mathrm{d}\sigma}\circ{\scriptstyle\cdots}\circ T_{a_{m}\mathrm{d}\sigma}$.
@@ 4 Asymptotic of the trace of a multiple composition
asymptotically as $\alpha\rightarrow\infty$ using stationary phase method through the parametrization $\Xi$. The stationary phase will be made over $\boldsymbol{s}$ with $t$ fixed, and then we will integrate over $t$. Regarding $\mathcal{V}$ as a subset of the product manifold $\Gamma^{m}$, its Riemannian metric $G_{\Xi}$ satisfies $\det G_{\Xi}(t,\boldsymbol{s})={\det G(t)}\cdot{\det G(t+s_{1})}\cdots{\det G(t+s_{m-1})}$. Make $\varphi(t,\boldsymbol{s})=\Phi\circ\Xi=\Phi\big{(}\gamma(t),\gamma(t+s_{1}),{\scriptstyle\dots},\gamma(t+s_{m-1})\big{)}$ so that we have
@@
We now use (4) to compute the Hessian of $\varphi$ at $\boldsymbol{s}=\boldsymbol{0}$. Let $11\leq q^{\prime}\leq m-$ 1 and $1\leq k\leq d$. Notice that if $q^{\prime}\notin\{q-1,q,q+1\}$, then (4) does not depend on $s^{(q^{\prime})}_{k}$, and therefore $\frac{\partial^{2}}{\partial s^{(q^{\prime})}_{k}\partial s^{(q)}_{j}}\varphi(t,\boldsymbol{s})=0$. For the remaining three no null cases we have
@@
In Appendix B it is shown that $\det\mathcal{O}_{m-1}$ is positive, while so is $\det G$ because $G$ is positive-definite. This implies that $\det\left(-i\mathrm{Hess}\varphi(t,\mathbf{0})\right)$ is also positive and so the application of the stationary phase method is justified. Apply Theorem 7.7.5 [5] to estimate the integral inside the brackets in (4), to obtain
@@
Substituting this on (4) we get
@@
Let $\gamma_{1},{\scriptstyle\dots},\gamma_{L}:B(0,1)\subset\mathbb{R}^{d}\rightarrow\Gamma$ be parametrizations of $\Gamma$ such that $\cup_{\ell=1}^{L}\gamma_{\ell}(B(0,1))=\Gamma$. Without loss of generality, we have taken this as a finite union since, even if $\Gamma$ is not compact, the integral in which we are interested is done on the supports of the functions $a_{j}$ which have compact support. Take $\varepsilon$ small enough and define
@@
is an open neighborhood of the diagonal in $\Gamma^{m}$ and $\left\{\Xi_{\ell},B(0,\varepsilon)^{m}\right\}_{\ell=1,\dots,L}$ is an atlas for $\mathcal{U}$. It follows from (15) and Lemma 3.3 that $\mathrm{Tr}(S)$ decays exponentially fast outside $\mathcal{U}$ when $\alpha$ grows. Let $\{\chi_{\ell}\}_{\ell=1,\dots,L}$ be a partition of unity subordinated to the aforementioned atlas. Applying what we obtained in (4) to each $\chi_{\ell}$ and substituting it into (15) we get
@@
In order to obtain an explicit formula from (4), we would like $\det\mathcal{O}_{m-1}$ to be computable, while it would be ideal for it to be constant for all $\xi=\gamma(t)\in\Gamma$. This does not need to be true in general, but in Appendix B we show that this is the case when $\Gamma$ is isotropic or co-isotropic. Namely, in Proposition B.11 we show that
@@
[Asymptotics for the trace of a composition] Let $\Gamma$ be isotropic or co-isotropic and $a_{1},{\scriptstyle\dots},a_{m}\in C^{2}_{0}(\Gamma)$. Then
@@
By the linearity of the trace, the result will follow from (4) by substituting the last identity for $c_{\alpha}$ and the expression for $2^{\frac{d(m-1)}{2}}(\mathrm{det}\mathcal{O}_{m-1})^{-\frac{1}{2}}$ given in (27). ∎
5 Section with the main theorem
@@
5.1 The Szegö limit theorem for polynomials
@@
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{2}_{0}(\Gamma)$. Then
@@
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{2}_{0}(\Gamma)$. There exists a constant $C>0$ such that
@@
Here we have used that $\lim\limits_{\alpha\rightarrow\infty}\alpha^{\frac{1}{\alpha}}=1$. Since the constant $C$ is independent of $\alpha$, the statement is proved. ∎
@@
[Szegö limit theorem for polynomials] Let $\Gamma$ be isotropic or co-isotropic, $a\in C^{2}_{0}(\Gamma)$, and $P$ be a polynomial without constant term. Then
@@ 5.2 The Szegö limit theorem
Let $\Gamma\subset\mathbb{B}_{n}$ be a submanifold and $a\in C^{2}_{0}(\Gamma)$ a positive function. For $p\in(0,1)$
@@
Let $\Gamma$ be isotropic or co-isotropic, $a\in C^{2}_{0}(\Gamma)$ a positive function, and $P$ a polynomial without constant term. For $p\in(0,1)$,
@@ 6 Corollaries and follow up results
In this section we present some results that follow as applications of the Szegö limit theorem 1.8. In particular, we obtain results describing the distribution of the eigenvalues and the Schatten norms of $\widehat{T}_{a\mathrm{d}\sigma}$.
@@ 6.1 Distribution of eigenvalues
$\lim\limits_{\alpha\rightarrow\infty}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}\left(\varphi\left(\widehat{T}_{a\mathrm{d}\sigma}\right)\right)=\frac{1}{2^{\frac{d^{\prime}}{2}}}\int_{0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\varphi(s)\mathcal{D}_{a}(s)\ \mathrm{d}s\ .$ (39)
@@
Proof.
Let $\mathbb{I}_{I}$ be the indicator function of $I$, which is continuous on each spectrum $\sigma(\widehat{T}_{a\mathrm{d}\sigma})$, $\alpha>0$. If $\{\lambda_{j}\}_{j\in\mathbb{N}}$ are the singular values of $\widehat{T}_{a\mathrm{d}\sigma}$ we notice that
$\mathcal{N}_{I}(\widehat{T}_{a\mathrm{d}\sigma})$ $=$ $\sum_{\{j\ :\ \lambda_{j}\in I\}}1\ =\ \sum_{j=1}^{\infty}\mathbb{I}_{I}(\lambda_{j})\ =\ \mathrm{Tr}\left(\mathbb{I}_{I}\left(\widehat{T}_{a\mathrm{d}\sigma}\right)\right)\ .$
Let $I=[c,d]$ and consider the two sequences of trapezoidal functions $\{\psi_{j}\}_{j\in\mathbb{N}}$, $\{\phi_{j}\}_{j\in\mathbb{N}}$ defined by
$\psi_{j}(x)=\left\{\begin{array}[]{cl}0&x\in[0,c]\\ j(x-c)&x\in(c,c+\frac{1}{j})\\ 1&x\in[c+\frac{1}{j},d-\frac{1}{j}]\\ -j(x-d)&x\in(d-\frac{1}{j},d)\\ 0&x\in[d,\infty)\end{array}\right.\quad,$
$\phi_{j}(x)=\left\{\begin{array}[]{cl}0&x\in[0,c-\frac{1}{j}]\\ j(x-c)+1&x\in(c-\frac{1}{j},c)\\ 1&x\in[c,d]\\ -j(x-d)+1&x\in(d,d+\frac{1}{j})\\ 0&x\in[d+\frac{1}{j},\infty)\end{array}\right.\ .$
These functions satisfy
$\lim\limits_{j\rightarrow\infty}\psi_{j}(x)=\mathbb{I}_{(c,d)}(x)\qquad,\qquad\lim\limits_{j\rightarrow\infty}\phi_{j}(x)=\mathbb{I}_{I}(x)$
for all $x>0$ and $\psi_{j}(x)\leq\mathbb{I}_{I}(x)\leq\phi_{j}(x)$ for all $j\in\mathbb{N}$ and $x>0$. It follows that
$\mathrm{Tr}\left(\psi_{j}(\widehat{T}_{a\mathrm{d}\sigma})\right)\ \leq\ \mathcal{N}_{I}(\widehat{T}_{a\mathrm{d}\sigma})\ \leq\ \mathrm{Tr}\left(\phi_{j}(\widehat{T}_{a\mathrm{d}\sigma})\right)\ ,$
and by (39)
$\int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\psi_{j}(s)\mathcal{D}_{a}(s)\ \mathrm{d}s\ \leq\ \frac{1}{2^{\frac{d^{\prime}}{2}}}\lim\limits_{\alpha\rightarrow\infty}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathcal{N}_{I}(\widehat{T}_{a\mathrm{d}\sigma})\\ \leq\ \int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\phi_{j}(s)\mathcal{D}_{a}(s)\ \mathrm{d}s$ (40)
for all $j\in\mathbb{N}$. Now note that $|\psi_{j}(x)|\leq|\phi_{j}(x)|\leq\phi_{1}(x)$ for all $x>0$, $j\in\mathbb{N}$ and $\int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\phi_{1}(s)\mathcal{D}_{a}(s)\ \mathrm{d}s<\infty$. Then by the dominated convergence theorem
$\lim\limits_{j\rightarrow\infty}\int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\psi_{j}(s)\mathcal{D}_{a}(s)\ \mathrm{d}s\ =\ \int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\mathbb{I}_{(c,d)}(x)\ \mathcal{D}_{a}(s)\ \mathrm{d}s$
and
$\lim\limits_{j\rightarrow\infty}\int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\phi_{j}(s)\mathcal{D}_{a}(s)\ \mathrm{d}s\ =\ \int_{\scriptscriptstyle 0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\mathbb{I}_{[c,d]}(x)\ \mathcal{D}_{a}(s)\ \mathrm{d}s\ ,$
both of them being equal to $\int_{I}\mathcal{D}_{a}(s)\ \mathrm{d}s$, which together with (40) concludes the proof. ∎
@@ 6.2 Schatten norm estimates
Let $p>0$. Recall that for $A$ a compact operator in $A^{2}_{\alpha}(\mathbb{B}_{n})$, $A$ is a member of the Schatten class $S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))$ if and only if $|A|^{p}$ is of trace class. This is equivalent to having the quantity

### v2

@@ 3.1 An integral expression for the composition
For the instance when $A=T_{a\mathrm{d}\sigma}$ is a Toeplitz operator, Fubini’s theorem and the reproducing property of the Bergman kernels yield in the expression
@@
We present a formula for the trace of a composition of $m$ Toeplitz operators
@@
Let $m\geq 2$ be a natural number, $a_{1},{\scriptstyle\dots},a_{m}\in L^{\infty}(\Gamma)$, and ${S:=T_{a_{1}\mathrm{d}\sigma}\circ{\scriptstyle\cdots}\circ T_{a_{m}\mathrm{d}\sigma}}$. Then
@@
Notice that $\frac{1-\langle\xi_{j},\xi_{j+1}\rangle}{1-|\xi_{j}|^{2}}\in(-\infty,0]$ if and only if $\langle\xi_{j},\xi_{j+1}\rangle\in[1,\infty)$, which does not occur since $\xi_{j},\xi_{j+1}\in\mathbb{B}_{n}$. It follows from (16) that $\Phi$ is well defined and continuous on $\Gamma^{m}$. Moreover, from (15) we find that all the $\xi_{j}$ involved in the integration are the ones in the support of $a_{j}$. If each $a_{j}$ has compact support in $\Gamma$, we conclude that there is an open neighborhood of $(-\infty,0]$ such that only values outside this neighborhood are used to evaluate $\operatorname{Log}$ in the expression for $\Phi$ when computing (15).
3.1.1 The behavior on the diagonal
We describe the behavior of $\mathrm{Im}\big{(}\Phi(\xi_{1},{\scriptstyle\dots},\xi_{m})\big{)}$ with the following result. From it we deduce that the critical behavior of $\Phi$ occurs on the diagonal of $\Gamma^{m}$.
@@ 3.1.1 The behavior on the diagonal
3.2 Asymptotic of the trace of a multiple composition
This section is devoted to finding the asymptotic behavior of the integral expression in (13) when $\alpha$ tends to infinity. Throughout the section, we will keep $m\geq 2$ and $a_{1},{\scriptstyle\dots},a_{m}\in C^{\infty}_{0}(\Gamma)$ fixed and denote $S:=T_{a_{1}\mathrm{d}\sigma}\circ{\scriptstyle\cdots}\circ T_{a_{m}\mathrm{d}\sigma}$.
@@ 3.2 Asymptotic of the trace of a multiple composition
asymptotically as $\alpha\rightarrow\infty$ using stationary phase method through the parametrization $\Xi(t,\boldsymbol{s})$. The stationary phase will be made over $\boldsymbol{s}$ with $t$ fixed and then we will integrate over $t$. Regarding $\mathcal{V}$ as a subset of the product manifold $\Gamma^{m}$, its Riemannian metric $G_{\Xi}$ satisfies $\det G_{\Xi}(t,\boldsymbol{s})={\det G(t)}\cdot{\det G(t+s_{1})}\cdots{\det G(t+s_{m-1})}$. Make $\varphi(t,\boldsymbol{s})=\Phi\circ\Xi=\Phi\big{(}\gamma(t),\gamma(t+s_{1}),{\scriptstyle\dots},\gamma(t+s_{m-1})\big{)}$ so that we have
@@
We now use (3.2) to compute the Hessian of $\varphi$ at $\boldsymbol{s}=\boldsymbol{0}$. Let $11\leq q^{\prime}\leq m-$ 1 and $1\leq k\leq d$. Notice that if $q^{\prime}\notin\{q-1,q,q+1\}$, then (3.2) does not depend on $s^{(q^{\prime})}_{k}$, and therefore $\frac{\partial^{2}}{\partial s^{(q^{\prime})}_{k}\partial s^{(q)}_{j}}\varphi(t,\boldsymbol{s})=0$. For the remaining three no null cases we have
@@
In Appendix B it is shown that $\det\mathcal{O}_{m-1}$ is positive, while so is $\det G$ because $G$ is positive-definite. This implies that $\det\left(-i\mathrm{Hess}\varphi(t,\mathbf{0})\right)$ is also positive and so the application of the stationary phase method is justified. Apply Theorem 7.7.5 [5] to estimate the integral inside the brackets in (3.2), to obtain
@@
Notice that up until this point we had not made use of the smoothness of the symbols $a_{1},{\scriptstyle\dots},a_{m}$. In order to apply the cited theorem to obtain the previous estimate, we need each $a_{j}\in C^{2k}_{0}(\Gamma)$ for some $k\geq\frac{d(m-1)}{2}+1$. As we wish for the hypotheses on the symbols to be independent of their quantity $m$, we have rather requested $a_{j}\in C^{\infty}_{0}(\Gamma)$ for each $j$. Also, as we will see later, this same $m$ will represent the exponent of powers of $\widehat{T}_{a\mathrm{d}\sigma}$. This is the reason why we requested $a\in C^{\infty}_{0}(\Gamma)$ in Definition 1.1.
Substituting on (3.2) the resulting estimate we get
@@
Let $\gamma_{1},{\scriptstyle\dots},\gamma_{L}:B(0,1)\subset\mathbb{R}^{d}\rightarrow\Gamma$ be parametrizations of $\Gamma$ such that $\cup_{\ell=1}^{L}\gamma_{\ell}(B(0,1))=\Gamma$. Without loss of generality, we have taken this as a finite union since, even if $\Gamma$ is not compact, the integral in which we are interested is done on the supports of the functions $a_{j}$. Take $\varepsilon$ small enough and define
@@
is an open neighborhood of the diagonal in $\Gamma^{m}$ and $\left\{\Xi_{\ell},B(0,\varepsilon)^{m}\right\}_{\ell=1,\dots,L}$ is an atlas for $\mathcal{U}$. It follows from (15) and Lemma 3.3 that $\mathrm{Tr}(S)$ decays exponentially fast outside $\mathcal{U}$ when $\alpha$ grows. Let $\{\chi_{\ell}\}_{\ell=1,\dots,L}$ be a partition of unity subordinated to the aforementioned atlas. Applying what we obtained in (3.2) to each $\chi_{\ell}$ and substituting it into (15) we get
@@
In order to obtain an explicit formula from (3.2), we would like $\det\mathcal{O}_{m-1}$ to be computable, while it would be ideal for it to be constant for all $\xi=\gamma(t)\in\Gamma$. This does not need to be true in general, but in Appendix B is shown this is the case when $\Gamma$ is isotropic or co-isotropic. Namely, Proposition B.10 states
@@
[Asymptotics for the trace of a composition] Let $\Gamma$ be isotropic or co-isotropic and $a_{1},{\scriptstyle\dots},a_{m}\in C^{\infty}_{0}(\Gamma)$. Then
@@
By the linearity of the trace, the result will follow from (3.2) by substituting the last identity for $c_{\alpha}$ and the expression for $2^{\frac{d(m-1)}{2}}(\mathrm{det}\mathcal{O}_{m-1})^{-\frac{1}{2}}$ given in (27). ∎
4 Proof of the main result
@@
4.1 The polynomial case
@@
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{\infty}_{0}(\Gamma)$. Then
@@ 4.1.1 An estimate for the operator norm
Let $\Gamma$ be isotropic or co-isotropic and $a\in C^{\infty}_{0}(\Gamma)$. There exists a constant $C>0$ such that
@@
Here we have used $\lim\limits_{\alpha\rightarrow\infty}\alpha^{\frac{1}{\alpha}}=1$. ∎
@@ 4.1.2 The Szegö limit theorem for polynomials
[Szegö limit theorem for polynomials] Let $\Gamma$ be isotropic or co-isotropic, $a\in C^{\infty}_{0}(\Gamma)$, and $P$ be a polynomial without constant term. Then
@@
Remark 4.6.
Notice that for this case the hypothesis regarding $a>0$ in Theorem 1.8 is not needed and is therefore omitted.
@@ 4.2 The Szegö limit theorem
Let $\Gamma\subset\mathbb{B}_{n}$ be a submanifold and $a\in C^{\infty}_{0}(\Gamma)$ a positive function. For $p\in(0,1)$
@@
Let $\Gamma$ be isotropic or co-isotropic, $a\in C^{\infty}_{0}(\Gamma)$ a positive function, and $P$ a polynomial without constant term. For $p\in(0,1)$,
@@ 5 Corollaries and follow up results
We present some applications of the Szegö limit theorem 1.8, concretely, results describing assymptotically the distribution of the eigenvalues and the Schatten norms of $\widehat{T}_{a\mathrm{d}\sigma}$. Both of these results are equivalent to the ones obtained in [11] and their proofs are analogues to the ones therein.
@@ 5.1 Distribution of eigenvalues
$\lim\limits_{\alpha\rightarrow\infty}\left(\frac{\pi}{\alpha}\right)^{\frac{d}{2}}\mathrm{Tr}\left(\varphi\left(\widehat{T}_{a\mathrm{d}\sigma}\right)\right)=\frac{1}{2^{\frac{d^{\prime}}{2}}}\int_{0}^{\scriptscriptstyle\max\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}}\varphi(s)\mathcal{D}_{a}(s)\ \mathrm{d}s\ .$
@@ 5.2 Schatten norm estimates
Let $p>0$. Recall that for $A$ a compact operator on $A^{2}_{\alpha}(\mathbb{B}_{n})$, $A$ is a member of the Schatten class $S_{p}(A^{2}_{\alpha}(\mathbb{B}_{n}))$ if and only if $|A|^{p}$ is of trace class. This is equivalent to having the quantity